jones87:

​Imogene's car traveled 392 mi averaging a certain speed. If the car had gone 7 mph​ faster, the trip would have taken 1 hour less. Find the average speed.

Moon:

speed = distance/time let's use x to represent the original speed and t to represent the original time x = 392/t "If the car had gone 7 mph​ faster, the trip would have taken 1 hour less." basically take the same 392 mile trip, add 7 to the speed and take away 1 from the hour to get: x + 7 = 392/(t-1) since we want x, solve the original equation for t so you can plug that into t for the second equation, re-writing the equation in terms of x only, and solve

Moon:

@smokeybrown can you check if I did this right I'm super tired/stressed

SmokeyBrown:

Yeah, that's right